OpenStudy (anonymous):

Divide.Assumed that all variables are nonzero,Its a polynomial question .dividing polynomial 35x^9y^10 ---------- 5x^8y^3

OpenStudy (anonymous):

I solved it and this is what I got 35x^9 divided by 5x^8=7x^9-8 and Y^10 divided by y^3=y^10-3

OpenStudy (anonymous):

7x^9-8 = 7x^1 and y^10-3=7x^7

OpenStudy (anonymous):

and my final answer was 7xy^7

OpenStudy (anonymous):

Do check for me if am right or not ...

OpenStudy (anonymous):

you're correct